# Tunneling in a Lorenz-like model for an active wave-particle entity

### v1

@@ VI Conclusions
The evolution of a normally distributed position density profile constructed from ensemble of initial conditions showed transmitted and reflected probability profiles after interacting with the potential barrier. Further, velocity oscillations and transient chaos, induced wave-like features in the transmitted and reflected probability density profiles after interacting with the barrier. Such wave-like features have been observed in transmitted and reflected density profiles of a 1D BEC interacting with a Gaussian potential barrier where both classical and quantum tunneling effects are present [35, 37].
In the context of walking droplets and hydrodynamic quantum analogs, although we used a simple model to explore the setup of a WPE tunneling across a potential barrier, the dynamical features of velocity oscillations, and also perhaps transient chaos, might be relevant for experiments. It would be interesting to revisit the experiments of a walking droplet tunneling by Eddi et al. [20] and Tadrist et al. [22], especially in the high memory regime where the free walker is undergoing velocity fluctuations before encountering the potential barrier. As demonstrated in this manuscript, this might result in further sensitivity to barrier crossing which would now be rooted in sensitivity to initial conditions due to unsteady dynamical features.

### v2

@@ VI Conclusions
The evolution of a normally distributed position density profile constructed from ensemble of initial conditions showed transmitted and reflected probability profiles after interacting with the potential barrier. Further, velocity oscillations and transient chaos, induced wave-like features in the transmitted and reflected probability density profiles after interacting with the barrier.
In the context of the experimental system of walking droplets, we note that our idealized model is not suitable for making quantitative predictions since it does not incorporate features of variations in vertical bouncing dynamics with memory [38], interactions with submerged barriers and wave modulation due to varying topography [39]. Moreover, the potential barrier in our model directly affects the droplet without affecting the droplet-generated waves. Conversely, in experiments of a walking droplet tunneling by Eddi et al. [21] and Tadrist et al. [23], the potential barrier is modeled as a submerged barrier whose effect on the droplet is mediated by the underlying waves. Hence, a more closer experimental setup to the model setup presented here would be to have a potential barrier in the form of an external magnetic field along with ferrofluid injected in the droplet, similar to the setup of Perrard et al. [40] created for generating a harmonic potential for the droplet. This method will ensure that the potential barrier is felt directly by the droplet without affecting the droplet-generated waves. Nevertheless, the key non-trivial qualitative features observed in this work, such as the stripes and intermingled patterns in the initial condition space and the enhanced sensitivity to barrier crossing, emerge from unsteady motion associated with velocity oscillations. Since the dynamical regime of a free walker undergoing velocity oscillations has been demonstrated experimentally [27], it would be interesting to revisit the experiments of a walking droplet tunneling by Eddi et al. [21] and Tadrist et al. [23] in the high memory regime, or perform experiments with a magnetic field potential barrier in the high memory regime, when the free walker is undergoing velocity fluctuations before encountering the barrier. As observed in our simple model, this might result in further sensitivity to barrier crossing which would now be rooted in sensitivity to initial conditions due to unsteady dynamical features. Lastly, in the context of generalized pilot-wave framework [9], our work demonstrates that nonlinear and chaotic features of a simple classical pilot-wave system provide a dynamical mechanism for the origin of sensitivity and unpredictability in barrier crossing.